OpenStudy (anonymous):

Why can't you divide by 0?

OpenStudy (amistre64):

becuase you cant get a result that makes sense

OpenStudy (amistre64):

10/5 = 2 , since 2*5 = 10 4/0 = ? , since ?*0 = 4 .....

OpenStudy (amistre64):

wheras: 0/3 = 0, since 0*3 = 0

OpenStudy (amistre64):

its simply undefined when we divide by 0 ...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

i think, it is because, no term other than we can get except 0 whenever we multiply something with 0

OpenStudy (anonymous):

so, there can be no quotient when a number is divided by 0